# SeaEval for Multilingual Foundation Models: From Cross-Lingual Alignment to Cultural Reasoning

### v3

@@ Abstract
[…]
@@ 1 Introduction
Language English Chinese Indonesian Spanish Question Please choose the correct answer for 回答下面问题,选择正确答案。 Silakan pilih jawaban yang benar Por favor elija la respuesta correcta the following question. 当白光通过棱镜时,比绿光弯曲 untuk pertanyaan berikut. para la siguiente pregunta. When white light passes through a 更多的光是? Ketika cahaya putih melewati sebuat Cuando la luz blanca pasa a través prism, the light that bends more than (A) 红色的 prisma, cahaya manakah yang de un prisma, la luz que se desvía green is? (B) 黄色的 memiliki sudut deviasi lebih besar más que la verde es (A) Red (C) 蓝色的 daripada cahaya hijau? (A) Rojo (B) Yellow (D) 都不是 (A) Merah (B) Amarillo (C) Blue (B) Kuning (C) Azul (D) None of these (C) Biru (D) Ninguna de las anteriores (D) Tak ada satupun Answer The correct answer is: (A) 红色的 (D) Tak ada satupun La respuesta correcta es: (C) Blue In English: (A) Red In English: (D) None of them (A) Rojo In English: (A) Red Correctness
@@ 2.1 Multilingual Foundation Model
Reasoning Capability Reasoning has long been treated as a complex yet essential capability in cognition that goes beyond fundamental language understanding McCarthy (2022), which can be demonstrated via extracting critical information from the text environment and drawing correct conclusions Nilsson (1991). The advancement of natural language reasoning has evolved from explicit and superficial reading comprehension and natural language inference Rajpurkar et al. (2016); Wang et al. (2019) to encompass implicit, complex, and specific reasoning capabilities such as multi-hop reasoning, numerical reasoning, and logical reasoning Lai et al. (2017); Huang et al. (2019); Jiao et al. (2022). While reasoning can be challenging due to the various relations and expressions involved, which are difficult to transcribe into symbolic or formal languages, FMs are shown to serve as a proxy to compress abundant knowledge, and solve various tasks following human instructions with less specialization Wei et al. (2022); Wang et al. (2023b); Jiao et al. (2023).
@@ Acknowledgement
This research was supported by the Institute for Infocomm Research (I2R), A*STAR, Singapore. The computational work for this article was partially performed on resources of the National Supercomputing Centre (NSCC), Singapore. It is also partially supported by Cloud TPUs from Google’s TPU Research Cloud (TRC). We thank Xunlong Zou for participating in research discussions, and Siti Umairah Md Salleh, Siti Maryam Binte Ahmad Subaidi, Nabilah Binte Md Johan, and Wiwik Karlina for their contribution to cross-lingual resource construction and verification.
@@ Appendix G Annotators
Multicultural and Multilingual Understanding SG-Eval Instruction Please carefully read the following question and select the most appropriate answer from the choices. Sample Which drink in Singapore has the highest calories? (A) Teh O (B) Teh Siew Dai (C) Kopi (D) Kopi C Answer: (C) Kopi US-Eval Instruction Read the following question carefully and select the correct answer from the choices. Sample When daylight-saving time arrives in the spring how do most Americans turn their clocks? (A) one hour forward (B) one hour backward (C) two hours forward (D) two hours backward Answer: (A) one hour forward CN-Eval Instruction 请仔细阅读以下问题,并从选项中选择最合适的答案。 Sample 清代官场饮茶有着特殊的程序和含义,有别于一般的茶道,主人若端茶,对客人说“请喝茶”,这表明 (A)对客人不满 (B)请客人品茶 (C)对客人的尊敬 (D)会谈结束送客 答案: (D)会谈结束送客 Translation: Tea drinking in officialdom in the Qing Dynasty had special procedures and meanings, which were different from ordinary tea ceremonies. If the host served tea and said "Please drink tea" to the guests, this meant (A) Dissatisfied with the guest (B) Invite guest to taste tea (C) Show Respect for guest (D) End the meeting and seeing off the guest Answer: (D) End the meeting and see off the guest Singlish2English Instruction Translate the following sentence from Singlish to English. Please only output the translated sentence. Sample Source in Singlish: Wah this one damn shiok and underrated. The maggi goreng also damn sedap. Bro you got refined taste Target in Standard English: Wow, this is super enjoyable and underrated. The Maggi Goreng is damn delicious. Brother, you have got a refined taste. Cross-Lingual Consistency Cross-MMLU Instruction Respond to the question by selecting the most appropriate answer. Sample Shown in Table 1. Cross-LogiQA Instruction Kindly choose the correct answer from the options provided for the multiple-choice question. Sample English Version: Content: At a gathering at which bankers, athletes, and lawyers are present, all of the bankers are athletes and none of the lawyers are bankers. Question: If the statements above are true, which one of the following statements must also be true? (A) Some of the lawyers are not athletes. (B) Some of the athletes are not lawyers. (C) None of the lawyers are athletes. (D) All of the athletes are bankers. Answer: (B) Some of the athletes are not lawyers. Chinese Version: 在银行家,运动员和律师的聚会上,所有银行家都是运动员,没有一个律师是银行家。 如果陈述以上为真,下列哪一项也一定为真? (A)有些律师不是运动员。 (B)有些运动员不是律师。 (C)没有律师是运动员。 (D)所有运动员都是银行家。 答案: (B)有些运动员不是律师。 Indonesian Version: Isi: Pada pertemuan yang dihadiri oleh para bankir, atlet, dan pengacara, semua bankir adalah atlet dan tidak ada satupun pengacara yang merupakan bankir. Pertanyaan: Jika pernyataan di atas benar, manakah pernyataan berikut yang juga benar? (A) Beberapa pengacara bukanlah atlet. (B) Beberapa atlet bukan pengacara. (C) Tidak ada pengacara yang merupakan atlet. (D) Semua atletnya adalah bankir. Jawaban: (B) Beberapa atlet bukan pengacara. Same sample in Spain, Vietnamese and Malay languages …
@@
Language Korean Malay Turkish German Question 다음 문제에서 올바른 답을 선택해 주세요. Sila pilih jawapan yang betul untuk soalan berikut. Lütfen aşağıdaki soru için doğru cevabı seçin. Bitte wählen Sie die richtige Antwort für die 흰 빛이 프리즘을 통과할 때, 초록보다 Apabila cahaya putih melalui prisma, cahaya yang Beyaz ışık bir prizmadan geçtiğinde yeşilden folgende Frage. Welches Licht wird stärker 더 많이 굴절되는 빛은 무엇입니까? membengkok lebih daripada hijau ialah daha fazla kırılan ışık gebeugt als grünes Licht wenn weißes Licht (A) 빨강 (A) Merah (A) kırmızıdır durch ein Prisma fällt? (B) 노랑 (B) Kuning (B) sarıdır (A) Rot (C) 파랑 (C) Biru (C) mavidir (B) Gelb (D) 이 중 어느 것도 아님 (D) Tiada satu pun daripada ini (D) Bunlardan hiçbiri (C) Blau (D) Keine von diesen Answer (C) 파랑 (A) Merah (C) mavidir (A) Rot In English: (C) Blue In English: (A) Red In English: (C) Blue In English: (A) Red Correctness

### v4

@@ Abstract
Bin Wang1,2121,21 , 2*, Zhengyuan Liu1,2121,21 , 2*, Xin Huang1111, Fangkai Jiao3333, Yang Ding1111, Ai Ti Aw1111, Nancy F. Chen1,3,41341,3,41 , 3 , 4 $1$Institute for Infocomm Research (I2R), A*STAR, Singapore $2$CNRS@CREATE, Singapore  $3$Nanyang Technological University, Singapore $4$Centre for Frontier AI Research (CFAR), A*STAR, Singapore {wang_bin,liu_zhengyuan,nfychen}@i2r.a-star.edu.sg
@@ 1 Introduction
Language English Chinese Indonesian Spanish Question Please choose the correct answer for {CJK*}UTF8gbsn 回答下面问题,选择正确答案。 Silakan pilih jawaban yang benar Por favor elija la respuesta correcta the following question. {CJK*}UTF8gbsn 当白光通过棱镜时,比绿光弯曲 untuk pertanyaan berikut. para la siguiente pregunta. When white light passes through a {CJK*}UTF8gbsn 更多的光是? Ketika cahaya putih melewati sebuat Cuando la luz blanca pasa a través prism, the light that bends more than {CJK*}UTF8gbsn (A) 红色的 prisma, cahaya manakah yang de un prisma, la luz que se desvía green is? {CJK*}UTF8gbsn (B) 黄色的 memiliki sudut deviasi lebih besar más que la verde es (A) Red {CJK*}UTF8gbsn (C) 蓝色的 daripada cahaya hijau? (A) Rojo (B) Yellow {CJK*}UTF8gbsn (D) 都不是 (A) Merah (B) Amarillo (C) Blue (B) Kuning (C) Azul (D) None of these (C) Biru (D) Ninguna de las anteriores (D) Tak ada satupun Answer The correct answer is: {CJK*}UTF8gbsn (A) 红色的 (D) Tak ada satupun La respuesta correcta es: (C) Blue In English: (A) Red In English: (D) None of them (A) Rojo In English: (A) Red Correctness
@@ 2.1 Multilingual Foundation Model
Reasoning Capability Reasoning has long been treated as a complex yet essential capability in cognition that goes beyond fundamental language understanding McCarthy (2022), which can be demonstrated via extracting critical information from the text environment and drawing correct conclusions Nilsson (1991). The advancement of natural language reasoning has evolved from explicit and superficial reading comprehension and natural language inference Rajpurkar et al. (2016); Wang et al. (2019); Wang and Li (2023) to encompass implicit, complex, and specific reasoning capabilities such as multi-hop reasoning, numerical reasoning, and logical reasoning Lai et al. (2017); Huang et al. (2019); Jiao et al. (2022). While reasoning can be challenging due to the various relations and expressions involved, which are difficult to transcribe into symbolic or formal languages, FMs are shown to serve as a proxy to compress abundant knowledge, and solve various tasks following human instructions with less specialization Wei et al. (2022); Wang et al. (2023b); Jiao et al. (2023).
@@ Acknowledgement
This work is supported by the National Research Foundation, Singapore under its AI Singapore Programme (AISG Award No: AISG2-GC-2022-005) and DesCartes: The National Research Foundation, Prime Minister’s Office, Singapore under its Campus for Research Excellence and Technological Enterprise (CREATE) program. Any opinions, findings, conclusions, or recommendations expressed in this material are those of the author(s) and do not reflect the views of the National Research Foundation, Singapore. The computational work for this article was partially performed on resources of the National Supercomputing Centre (NSCC), Singapore. It is also partially supported by Cloud TPUs from Google’s TPU Research Cloud (TRC). We thank Xunlong Zou and Geyu Lin for participating in research discussions, and Siti Umairah Md Salleh, Siti Maryam Binte Ahmad Subaidi, Nabilah Binte Md Johan, Wiwik Karlina, Xuan Long Do, Fabian Ritter Gutierrez and Ayrton San Joaquin for their contribution to cross-lingual resource construction and verification.
@@ Appendix G Annotators
Multicultural and Multilingual Understanding SG-Eval Instruction Please carefully read the following question and select the most appropriate answer from the choices. Sample Which drink in Singapore has the highest calories? (A) Teh O (B) Teh Siew Dai (C) Kopi (D) Kopi C Answer: (C) Kopi US-Eval Instruction Read the following question carefully and select the correct answer from the choices. Sample When daylight-saving time arrives in the spring how do most Americans turn their clocks? (A) one hour forward (B) one hour backward (C) two hours forward (D) two hours backward Answer: (A) one hour forward CN-Eval Instruction {CJK*}UTF8gbsn 请仔细阅读以下问题,并从选项中选择最合适的答案。 Sample {CJK*}UTF8gbsn 清代官场饮茶有着特殊的程序和含义,有别于一般的茶道,主人若端茶,对客人说“请喝茶”,这表明 {CJK*}UTF8gbsn (A)对客人不满 {CJK*}UTF8gbsn (B)请客人品茶 {CJK*}UTF8gbsn (C)对客人的尊敬 {CJK*}UTF8gbsn (D)会谈结束送客 {CJK*}UTF8gbsn 答案: (D)会谈结束送客 Translation: Tea drinking in officialdom in the Qing Dynasty had special procedures and meanings, which were different from ordinary tea ceremonies. If the host served tea and said "Please drink tea" to the guests, this meant (A) Dissatisfied with the guest (B) Invite guest to taste tea (C) Show Respect for guest (D) End the meeting and seeing off the guest Answer: (D) End the meeting and see off the guest Singlish2English Instruction Translate the following sentence from Singlish to English. Please only output the translated sentence. Sample Source in Singlish: Wah this one damn shiok and underrated. The maggi goreng also damn sedap. Bro you got refined taste Target in Standard English: Wow, this is super enjoyable and underrated. The Maggi Goreng is damn delicious. Brother, you have got a refined taste. Cross-Lingual Consistency Cross-MMLU Instruction Respond to the question by selecting the most appropriate answer. Sample Shown in Table 1. Cross-LogiQA Instruction Kindly choose the correct answer from the options provided for the multiple-choice question. Sample English Version: Content: At a gathering at which bankers, athletes, and lawyers are present, all of the bankers are athletes and none of the lawyers are bankers. Question: If the statements above are true, which one of the following statements must also be true? (A) Some of the lawyers are not athletes. (B) Some of the athletes are not lawyers. (C) None of the lawyers are athletes. (D) All of the athletes are bankers. Answer: (B) Some of the athletes are not lawyers. Chinese Version: {CJK*}UTF8gbsn 在银行家,运动员和律师的聚会上,所有银行家都是运动员,没有一个律师是银行家。 {CJK*}UTF8gbsn 如果陈述以上为真,下列哪一项也一定为真? {CJK*}UTF8gbsn (A)有些律师不是运动员。 {CJK*}UTF8gbsn (B)有些运动员不是律师。 {CJK*}UTF8gbsn (C)没有律师是运动员。 {CJK*}UTF8gbsn (D)所有运动员都是银行家。 {CJK*}UTF8gbsn 答案: (B)有些运动员不是律师。 Indonesian Version: Isi: Pada pertemuan yang dihadiri oleh para bankir, atlet, dan pengacara, semua bankir adalah atlet dan tidak ada satupun pengacara yang merupakan bankir. Pertanyaan: Jika pernyataan di atas benar, manakah pernyataan berikut yang juga benar? (A) Beberapa pengacara bukanlah atlet. (B) Beberapa atlet bukan pengacara. (C) Tidak ada pengacara yang merupakan atlet. (D) Semua atletnya adalah bankir. Jawaban: (B) Beberapa atlet bukan pengacara. Same sample in Spain, Vietnamese and Malay languages …
@@
Language Korean Malay Turkish German Question {CJK}UTF8mj다음 문제에서 올바른 답을 선택해 주세요. Sila pilih jawapan yang betul untuk soalan berikut. Lütfen aşağıdaki soru için doğru cevabı seçin. Bitte wählen Sie die richtige Antwort für die {CJK}UTF8mj흰 빛이 프리즘을 통과할 때, 초록보다 Apabila cahaya putih melalui prisma, cahaya yang Beyaz ışık bir prizmadan geçtiğinde yeşilden folgende Frage. Welches Licht wird stärker {CJK}UTF8mj더 많이 굴절되는 빛은 무엇입니까? membengkok lebih daripada hijau ialah daha fazla kırılan ışık gebeugt als grünes Licht wenn weißes Licht {CJK}UTF8mj(A) 빨강 (A) Merah (A) kırmızıdır durch ein Prisma fällt? {CJK}UTF8mj(B) 노랑 (B) Kuning (B) sarıdır (A) Rot {CJK}UTF8mj(C) 파랑 (C) Biru (C) mavidir (B) Gelb {CJK}UTF8mj(D) 이 중 어느 것도 아님 (D) Tiada satu pun daripada ini (D) Bunlardan hiçbiri (C) Blau (D) Keine von diesen Answer {CJK}UTF8mj(C) 파랑 (A) Merah (C) mavidir (A) Rot In English: (C) Blue In English: (A) Red In English: (C) Blue In English: (A) Red Correctness